About strikes anunder water rock and opens a pencil-sized crack 7.0mmwide 

min to locate the crack and plug it up. How much water has entered the boat in the meantime?


1
Expert's answer
2021-08-10T17:57:18-0400

Consider great area to be length"\times" width

="150\times7=10.5cm^2"

Now the depth of crack below water line=65cm

Newton's 2nd law

"v^2=u^2+2as"

u=0

"v^2=2gs"


"v^2=2\times980\times65=127400\\v=357 cm/sec"

Flow rate Q=Av

"Q=357\times10.5=3747cm^3/sec\\Q=3.74lit"

t=5min =300sec

Total volume ="3.74\times300=1122titer=1.12m^3"